# Gradient Cuff: Detecting Jailbreak Attacks on Large Language Models by Exploring Refusal Loss Landscapes

### v2

@@ Abstract
Large Language Models, Jailbreak Detection, AI Alignment and Safety
@@ 1 Introduction
Figure 1: Overview of Gradient Cuff. (a) introduces an example of jailbreak prompts by presenting a conversation between malicious actors and the Vicuna chatbot. (b) visualizes the refusal loss landscape for malicious queries and benign queries by plotting the interpolation of two random directions in the query embedding with coefficients $\alpha$ and $\beta$ following (Li et al., 2018). The refusal loss evaluates the probability that the LLM would not directly reject the input query, and the loss value is computed using Equation 3. See details of how to plot (b) in Appendix A.4. (c) shows the running flow of Gradient Cuff (at top), practical computing examples for refusal loss (at bottom left), and the distributional difference of the gradient norm of refusal loss on benign and malicious queries (bottom right). (d) shows the performance of Gradient Cuff against 6 jailbreak attacks for Vicuna-7B-V1.5. See Appendix A.6 for full results.
With the stupendous success of large language models (LLMs) such as GPT-4 (OpenAI, 2023), LLaMA-2 (Touvron et al., 2023), and Vicuna (Zheng et al., 2023), there is a trend to integrate these LLMs into various applications such as ChatGPT and Bing Search. In these applications, LLMs are used as the service backend. The front end of these applications receives the user input query from the interface, encapsulates it into a system prompt, and then sends it to the LLM to get a response. With the rapidly increasing social impact of these applications, model alignment and safety assurance to reduce harm and misuse have become significant considerations when developing and deploying LLMs. Methods such as Reinforcement Learning from Human Feedback (RLHF) have been proven to be an effective training technique to align LLMs with human values (Askell et al., 2021; Bai et al., 2022; Kasirzadeh & Gabriel, 2022; Ouyang et al., 2022).
@@
Many red-teaming efforts (Zou et al., 2023; Liu et al., 2023; Chao et al., 2023; Mehrotra et al., 2023; Wei et al., 2023a; Yong et al., 2023) have been put into designing algorithms to automatically generate jailbreak prompts to help test the robustness of aligned LLMs.
Specifically, GCG (Zou et al., 2023), one of the earlier works in this area, can successfully jailbreak several LLMs by optimizing an inserted universal adversarial suffix, which means that the embedded alignment effort is completely broken by the jailbreak attack.
Since the discovery of jailbreak risks for LLMs, various methods have been explored to defend against jailbreak attacks (Jain et al., 2023; Robey et al., 2023; Xie et al., 2023; Kumar et al., 2023) and have gained some success in detecting certain types of attacks such as GCG (Liu et al., 2023; Jain et al., 2023). However, in our systematic analysis, existing defenses either fail to be resistant against all types of jailbreak attacks, or have a significant detrimental effect on benign queries. PPL (Jain et al., 2023) uses an LLM to compute the perplexity of the input user query and filters those with perplexity greater than the threshold. PPL has been proven to have a good detection performance for GCG attacks but does not perform well on jailbreak prompts with good meaningfulness and fluency (Liu et al., 2023). Erase-Check (Kumar et al., 2023) and Self-Reminder (Xie et al., 2023) behave well on malicious queries but will misclassify many benign user queries, making the defenses overly conservative and impractical.
To alleviate the threats of jailbreak attacks and avoid the aforementioned problems in existing defenses, we propose Gradient Cuff, which detects jailbreak prompts by checking the refusal loss of the input user query and estimating the gradient norm of the loss function. We begin by introducing the concept of refusal loss and showcase the different behaviors of the loss function for benign instructions and malicious instructions.
A plot of the refusal loss landscape for benign and malicious instructions can be found in Figure 1 (b). By exploring the landscapes of refusal loss, we find that the refusal loss function for malicious instructions tends to have a smaller value and a larger gradient norm.
We then leverage this unique loss landscape characteristic to propose a two-step jailbreak detection algorithm, which is illustrated in Figure 1 (c). Figure 1 (d) evaluates 6 jailbreak attacks on Vicuna-7B-V1.5 and shows that our defense can reduce the attack success rate (ASR) averaged over these 6 jailbreaks from $74.3\%$ to $24.4\%$ on average. We also compare Gradient Cuff to other existing defense methods on Vicuna-7B-V1.5 and LLaMA-2-7B-Chat against these 6 jailbreaks as well as adaptive attacks to demonstrate our defense capabilities.
@@
  • We also show that Gradient Cuff is complementary to prompt-engineering based alignment strategies. When combined with Gradient Cuff, the performance of Self-Reminder, a system prompt design method (Xie et al., 2023), can be increased by a large margin.
@@ 2 Related Work
Jailbreak Attacks. Existing jailbreaks can be roughly divided into feedback-based jailbreak attacks and rule-based jailbreak attacks. Feedback-based jailbreaks utilize the feedback from the target LLM to iteratively update the jailbreak prompt until the model complies with the malicious instruction embedded in the jailbreak prompt. Feedback-based jailbreaks can be further categorized by their access mode to the target LLM. Some feedback-based jailbreak attacks like GCG (Zou et al., 2023), require white-box access to the target LLM. Specifically, GCG leverages gradients with respect to the one-hot token indicators to find better token choices at each position. Some feedback-based jailbreaks need gray-box access to the target LLM. The typical one is AutoDAN (Liu et al., 2023), which employs the target LLM’s generative loss of the target response to design the fitness score of the candidate jailbreak prompt to guide further optimization. PAIR  (Chao et al., 2023) and TAP (Mehrotra et al., 2023) are the representatives of feedback-based jailbreaks which only require black-box access to the target LLM. In PAIR and TAP, there are also two LLMs taking on the attacker role and the evaluator role. At each iteration, the attacker-generated jailbreak prompt would be rated and commented on by the evaluator model according to the target LLM’s response to the attack. Next, the attacker would generate new jailbreak prompts based on the evaluator’s comments, and repeat the above cycle until the jailbreak prompt can get full marks from the evaluator. The only information provided by the target LLM is the response to the jailbreak attack. As for the rule-based jailbreak attacks, we highlight Base64 (Wei et al., 2023a) and Low Resource Language (LRL) (Yong et al., 2023). Base64 encodes the malicious instruction into base64 format and LRL translates the malicious instruction into the language that is rarely used in the training process of the target LLM, such as German, Swedish, French and Chinese.
@@
Jailbreak Defenses. PPL (Jain et al., 2023) uses an LLM to compute the perplexity of the input query and rejects those with high perplexity. SmoothLLM (Robey et al., 2023), motivated by randomized smoothing (Cohen et al., 2019), perturbs the original input query to obtain several copies and aggregates the intermediate responses of the target LLM to these perturbed queries to give the final response to the original query. Erase-Check employs a model to check whether the original query or any of its erased subsentences is harmful. The query would be rejected if the query or one of its sub-sentences is regarded as harmful by the safety checker. Unlike these methods, there is another line of work (Xie et al., 2023; Zhang et al., 2024; Wei et al., 2023b; Varshney et al., 2024) where prompt engineering techniques are used to defend against jailbreak attacks. Notably, Self-Reminder (Xie et al., 2023) shows promising results by modifying the system prompt of the target LLM so that the model reminds itself to process and respond to the user in the context of being an aligned LLM.
@@ 3.1 Refusal Loss Function and Landscape Exploration
Current transformer-based LLMs will return different responses to the same query due to the randomness of autoregressive sampling based generation (Fan et al., 2018; Holtzman et al., 2020). With this randomness, it is an interesting phenomenon that a malicious user query will sometimes be rejected by the target LLM, but sometimes be able to bypass the safety guardrail. Based on this observation, for a given LLM $T_{\theta}$ parameterized with $\theta$, we define the refusal loss function $\phi_{\theta}(x)$ for a given input user query $x$ as below:
@@
where $y$ represents the response of $T_{\theta}$ to the input user query $x$. $JB(\cdot)$ is a binary indicator function to determine whether the response triggers a refusal action by the LLM. The function $p_{\theta}$ can be interpreted as the expected rate of getting refusal on the response $y$ from $T_{\theta}$ taking into account the randomness in the decoding process. Therefore, by our definition, the refusal loss function $\phi_{\theta}(x)$ can be interpreted as the likelihood of generating a non-refusal response to $x$. Following SmoothLLM (Robey et al., 2023), we define $JB(\cdot)$ as
@@
For example, $JB(y)$ would be 00 if $y=$”Sure, here is the python code to …” and $JB(y)$ would be $1$ if $y=$”Sorry, I cannot fulfill your request…”. We discuss more details about the implementation of the indicator function in Appendix A.3.
@@
$\phi_{\theta}(x)=~{}1-\mathbb{E}(Y)=1-p_{\theta}(x).\$
@@
where $\{Y_{i}|i=1,2,...,N\}$ is obtained by running $N$ independent realizations of the random variable $Y$. In the $i^{th}$ trial, we query the LLM $T_{\theta}$ using $x$ to get the response $y_{i}\sim T_{\theta}(x)$, and apply the indicator function $JB(\cdot)$ on $y_{i}$ to get $Y_{i}=JB(y_{i})$. Equation (3) can be explained as using the sample mean of the random variable $Y$ to approximate its expected value $\mathbb{E}(Y)$.
In general, $\phi_{\theta}(x)<0.5$ could be used as a naive detector to reject $x$ since $p_{\theta}(x)$ can be interpreted as the probability that $T_{\theta}$ regards $x$ as harmful. To further explore how this refusal loss can be used to improve jailbreak detection, we visualize the refusal loss landscape following the 2-D visualization techniques from (Li et al., 2018) in Figure 1 (b). From Figure 1 (b), we find that the landscape of $f_{\theta}(\cdot)$ is more precipitous for malicious queries than for benign queries, which implies that $f_{\theta}(\cdot)$ tends to have a large gradient norm if $x$ represents a malicious query. This observation motivates our proposal of using the gradient norm of $f_{\theta}(\cdot)$ to detect jailbreak attempts that pass the initial filtering of rejecting $x$ when $f_{\theta}(x)<0.5$.
@@ 3.2 Gradient Norm Estimation
In general, the exact gradient of $\phi_{\theta}(x)$ (or $f_{\theta}(x)$) is infeasible to obtain due to the existence of discrete operations such as applying the $JB(\cdot)$ function to the generated response, and the possible involvement of black-box evaluation functions (e.g., Perspective API). We propose to use zeroth order gradient estimation to compute the approximate gradient norm, which is widely used in black-box optimization with only function evaluations (zeroth order information) (Berahas et al., 2022; Liu et al., 2020). Similar gradient estimation techniques were used to generate adversarial examples from black-box models (Chen et al., 2017; Ilyas et al., 2018; Cheng et al., 2019).
@@
Based on this definition, we approximate the exact gradient $\nabla\phi_{\theta}(x)$ by $g_{\theta}(x)$, which is the estimated gradient of $f_{\theta}(x)$. Following (Berahas et al., 2022; Liu et al., 2020), we calculate $g_{\theta}(x)$ using the directional derivative approximation
@@
This theorem demonstrates that one can reduce the approximation error by taking larger values for $N$ and $P$. We provide the proof in Appendix A.10.
@@ 3.3 Gradient Cuff: Two-step jailbreak detection
With the discussions in Section 3.1 and Section 3.2, we now formally propose Gradient Cuff, a two-step jailbreak detection method based on checking the refusal loss and its gradient norm. Our detection procedure is shown in Figure 1 (c). Gradient Cuff can be summarized into two phases:
  • (Phase 1) Sampling-based Rejection: In the first step, we reject the user query $x$ by checking whether $f_{\theta}(x)<0.5$. If true, then $x$ is rejected, otherwise, $x$ is pushed into phase 2.
  • (Phase 2) Gradient Norm Rejection: In the second step, we regard $x$ as having jailbreak attempts if the norm of the estimated gradient $g_{\theta}(x)$ is larger than a configurable threshold $t$, i.e., $\|g_{\theta}(x)\|>t$.
@@
We summarize our method in Algorithm 1. The algorithm is implemented by querying the LLM $T_{\theta}$ multiple times, each to generate a response for the same input query $x$. Let $t_{\theta}(x)$ be the time cost for $T_{\theta}$ to generate a response for $x$. The dominating time cost $C_{\theta}$ for Gradient Cuff is
$C_{\theta}(N,P,x)=N\cdot(P+1)\cdot t_{\theta}(x),$ (7)
where $q=N\cdot(P+1)$ is the total query times to $T_{\theta}$ required to compute $f_{\theta}(x)$ and $g_{\theta}(x)$ in Gradient Cuff.
Algorithm 1 Gradient Cuff: Two-Step Jailbreak Detection
  Notations
1:  The LLM to be protected: $T_{\theta}$
2:  Required benign refusal (false positive) rate: $\sigma$
3:  Gaussian vector numbers: $P$
4:  LLM Response Sampling numbers: $N$
5:  Smoothing parameter: $\mu$
6:  Collection of benign user queries: $\mathcal{B}_{val}$
7:  Threshold: $t$
8:  Input User Query: $x_{test}$
  Threshold Selection
[…]
2:  Construct $G=\{\|g_{\theta}(x)\|~{}|~{}~{}x\in\mathcal{B}_{val}\setminus S\}$ based on equation (6).
3:  Sort $G$ in descending order
4:  Select $k$ that fulfills: $k-1\leq|\mathcal{B}_{val}|\cdot\sigma-|S|<k$
5:  Set threshold $t=G[k]$  # such that $\frac{|S|+k-1}{|\mathcal{B}_{val}|}\leq\sigma$
  Detection on test query $x_{test}$
1:  Calculate $f_{\theta}(x_{test})$ based on Equation (3).
2:  if $f_{\theta}(x_{test})<0.5$ then
3:      return ”I cannot fulfill your request.”
4:      # First-step filtering
5:  else
6:      Calculate $g_{\theta}(x_{test})$ based on Equation (6).
7:      if $g_{\theta}(x_{test})>t$ then
8:          return ”I cannot fulfill your request.”
9:          # Second-step filtering
10:       else
11:          return $y\sim T_{\theta}(x_{test})$  # No rejection

### v3

@@ 1 Introduction
Figure 1: Overview of Gradient Cuff. (a) introduces an example of jailbreak prompts by presenting a conversation between malicious actors and the Vicuna chatbot. (b) visualizes the refusal loss landscape for malicious queries and benign queries by plotting the interpolation of two random directions in the query embedding with coefficients $\alpha$ and $\beta$ following [15]. The refusal loss evaluates the probability that the LLM would not directly reject the input query, and the loss value is computed using Equation 3. See details of how to plot (b) in Appendix A.4. (c) shows the running flow of Gradient Cuff (at top), practical computing examples for refusal loss (at bottom left), and the distributional difference of the gradient norm of refusal loss on benign and malicious queries (bottom right). (d) shows the performance of Gradient Cuff against 6 jailbreak attacks for Vicuna-7B-V1.5. See Appendix A.6 for full results.
With the stupendous success of large language models (LLMs) such as GPT-4 [19], LLaMA-2 [23], and Vicuna [32], there is a trend to integrate these LLMs into various applications such as ChatGPT and Bing Search. In these applications, LLMs are used as the service backend. The front end of these applications receives the user input query from the interface, encapsulates it into a system prompt, and then sends it to the LLM to get a response. With the rapidly increasing social impact of these applications, model alignment and safety assurance to reduce harm and misuse have become significant considerations when developing and deploying LLMs. Methods such as Reinforcement Learning from Human Feedback (RLHF) have been proven to be an effective training technique to align LLMs with human values [1, 2, 13, 20].
@@
Many red-teaming efforts [33, 17, 4, 18, 26, 30] have been put into designing algorithms to automatically generate jailbreak prompts to help test the robustness of aligned LLMs. Specifically, GCG [33], one of the earlier works in this area, can successfully jailbreak several LLMs by optimizing an inserted universal adversarial suffix. This finding suggests that the embedded alignment effort in LLMs could be completely broken by the jailbreak attack.
Since the discovery of jailbreak risks for LLMs, various methods have been explored to defend against jailbreak attacks [12, 21, 28, 14] and have gained some success in detecting certain types of attacks such as GCG [17, 12]. However, in our systematic analysis, existing defenses either fail to be resistant against all types of jailbreak attacks, or have a significant detrimental effect on benign queries. PPL [12] uses an LLM to compute the perplexity of the input user query and filters those with perplexity greater than the threshold. PPL has been proven to have a good detection performance for GCG attacks but does not perform well on jailbreak prompts with good meaningfulness and fluency [17]. Erase-Check [14] and Self-Reminder [28] behave well on malicious queries but will misclassify many benign user queries, making the defenses overly conservative and impractical.
To alleviate the threats of jailbreak attacks and avoid the aforementioned problems in existing defenses, we propose Gradient Cuff, which detects jailbreak prompts by checking the refusal loss of the input user query and estimating the gradient norm of the loss function. We begin by introducing the concept of refusal loss and showcase the different behaviors of the loss function for benign instructions and malicious instructions. A plot of the refusal loss landscape for benign and malicious instructions can be found in Figure 1 (b). By exploring the landscapes of refusal loss, we find that the refusal loss function for malicious instructions tends to have a smaller value and a larger gradient norm.
We then leverage this unique loss landscape characteristic to propose a two-step jailbreak detection algorithm, which is illustrated in Figure 1 (c). Figure 1 (d) evaluates 6 jailbreak attacks on Vicuna-7B-V1.5 and shows that our defense can reduce the attack success rate (ASR) averaged over these 6 jailbreaks from $76.7\%$ to $25.7\%$ on average. We also compare Gradient Cuff to other existing defense methods on Vicuna-7B-V1.5 and LLaMA-2-7B-Chat against these 6 jailbreaks as well as adaptive attacks to demonstrate our defense capabilities.
@@
  • We also show that Gradient Cuff is complementary to prompt-engineering based alignment strategies. When combined with Gradient Cuff, the performance of Self-Reminder, a system prompt design method [28], can be increased by a large margin.
@@ 2 Related Work
Jailbreak Attacks. Existing jailbreaks can be roughly divided into feedback-based jailbreak attacks and rule-based jailbreak attacks. Feedback-based jailbreaks utilize the feedback from the target LLM to iteratively update the jailbreak prompt until the model complies with the malicious instruction embedded in the jailbreak prompt. Feedback-based jailbreaks can be further categorized by their access mode to the target LLM. Some feedback-based jailbreak attacks like GCG [33], require white-box access to the target LLM. Specifically, GCG leverages gradients with respect to the one-hot token indicators to find better token choices at each position. Some feedback-based jailbreaks need gray-box access to the target LLM. The typical one is AutoDAN [17], which employs the target LLM’s generative loss of the target response to design the fitness score of the candidate jailbreak prompt to guide further optimization. PAIR  [4] and TAP [18] are the representatives of feedback-based jailbreaks which only require black-box access to the target LLM. In PAIR and TAP, there are also two LLMs taking on the attacker role and the evaluator role. At each iteration, the attacker-generated jailbreak prompt would be rated and commented on by the evaluator model according to the target LLM’s response to the attack. Next, the attacker would generate new jailbreak prompts based on the evaluator’s comments, and repeat the above cycle until the jailbreak prompt can get full marks from the evaluator. The only information provided by the target LLM is the response to the jailbreak attack. As for the rule-based jailbreak attacks, we highlight Base64 [26] and Low Resource Language (LRL) [30]. Base64 encodes the malicious instruction into base64 format and LRL translates the malicious instruction into the language that is rarely used in the training process of the target LLM, such as German, Swedish, French and Chinese.
@@
Jailbreak Defenses. PPL [12] uses an LLM to compute the perplexity of the input query and rejects those with high perplexity. SmoothLLM [21], motivated by randomized smoothing [7], perturbs the original input query to obtain several copies and aggregates the intermediate responses of the target LLM to these perturbed queries to give the final response to the original query. Erase-Check employs a model to check whether the original query or any of its erased subsentences is harmful. The query would be rejected if the query or one of its sub-sentences is regarded as harmful by the safety checker. Another line of work [28, 31, 27, 25] use prompt engineering techniques to defend against jailbreak attacks. Notably, Self-Reminder [28] shows promising results by modifying the system prompt of the target LLM so that the model reminds itself to process and respond to the user in the context of being an aligned LLM. Unlike these unsupervised methods, some works like LLaMA-Guard [22] and Safe-Decoding [29] need to train an extra LLM. LLaMA-Guard trained a LLaMA-based model to determine whether the user query or model response contains unsafe content. Safe-Decoding finetuned the protected LLM on (malicious query, model refusal) pairs to get an expert LLM, and utilized the expert LLM to guide the safety-aware decoding during inference time. We will mainly focus on unsupervised methods as it is training-free and only need to deploy the protected LLM itself during inference time.
@@ 3.1 Refusal Loss Function and Landscape Exploration
Current transformer-based LLMs will return different responses to the same query due to the randomness of autoregressive sampling based generation [8, 10]. With this randomness, it is an interesting phenomenon that a malicious user query will sometimes be rejected by the target LLM, but sometimes be able to bypass the safety guardrail. Based on this observation, for a given LLM $T_{\theta}$ parameterized with $\theta$, we define the refusal loss function $\phi_{\theta}(x)$ for a given input user query $x$ as below:
@@
where $y$ represents the response of $T_{\theta}$ to the input user query $x$. $JB(\cdot)$ is a binary indicator function to determine whether the response triggers a refusal action by the LLM. The function $p_{\theta}$ can be interpreted as the expected rate of getting refusal on the response $y$ from $T_{\theta}$ taking into account the randomness in the decoding process. Therefore, by our definition, the refusal loss function $\phi_{\theta}(x)$ can be interpreted as the likelihood of generating a non-refusal response to $x$. Following SmoothLLM [21], we define $JB(\cdot)$ as
@@
For example, $JB(y)$ would be 00 if $y=$”Sure, here is the python code to …” and $JB(y)$ would be $1$ if $y=$"Sorry, I cannot fulfill your request…". We discuss more details about the implementation of the indicator function in Appendix A.3.
@@
$\phi_{\theta}(x)=~{}1-\mathbb{E}[Y]=1-p_{\theta}(x).\$
@@
where $\{Y_{i}|i=1,2,...,N\}$ is obtained by running $N$ independent realizations of the random variable $Y$. In the $i^{th}$ trial, we query the LLM $T_{\theta}$ using $x$ to get the response $y_{i}\sim T_{\theta}(x)$, and apply the indicator function $JB(\cdot)$ on $y_{i}$ to get $Y_{i}=JB(y_{i})$. Equation (3) can be explained as using the sample mean of the random variable $Y$ to approximate its expected value $\mathbb{E}[Y]$.
In general, $\phi_{\theta}(x)<0.5$ could be used as a naive detector to reject $x$ since $p_{\theta}(x)$ can be interpreted as the probability that $T_{\theta}$ regards $x$ as harmful. However, this detector alone only has limited effect against jailbreak attacks, as discussed in Section 4.3. To further explore how this refusal loss can be used to improve jailbreak detection, we visualize the refusal loss landscape following the 2-D visualization techniques from [15] in Figure 1 (b). From Figure 1 (b), we find that the landscape of $f_{\theta}(\cdot)$ is more precipitous for malicious queries than for benign queries, which implies that $f_{\theta}(\cdot)$ tends to have a large gradient norm if $x$ represents a malicious query. This observation motivates our proposal of using the gradient norm of $f_{\theta}(\cdot)$ to detect jailbreak attempts that pass the initial filtering of rejecting $x$ when $f_{\theta}(x)<0.5$.
@@ 3.2 Gradient Norm Estimation
In general, the exact gradient of $\phi_{\theta}(x)$ (or $f_{\theta}(x)$) is infeasible to obtain due to the existence of discrete operations such as applying the $JB(\cdot)$ function to the generated response, and the possible involvement of black-box evaluation functions (e.g., Perspective API).
We propose to use zeroth order gradient estimation to compute the approximate gradient norm, which is widely used in black-box optimization with only function evaluations (zeroth order information) [3, 16]. Similar gradient estimation techniques were used to generate adversarial examples from black-box models [5, 11, 6].
@@
Based on this definition, we approximate the exact gradient $\nabla\phi_{\theta}(x)$ by $g_{\theta}(x)$, which is the estimated gradient of $f_{\theta}(x)$. Following [3, 16], we calculate $g_{\theta}(x)$ using the directional derivative approximation
@@
This theorem demonstrates that one can reduce the approximation error by taking larger values for $N$ and $P$. We provide the proof in Appendix A.12. Experimental results in Appendix A.14 and Appendix A.13 also provide empirical evidence to support this theorem by demonstrating the scaling performance of Gradient Cuff with increased total queries.
@@ 3.3 Gradient Cuff: Two-step jailbreak detection
With the discussions in Section 3.1 and Section 3.2, we now formally propose Gradient Cuff, a two-step jailbreak detection method based on checking the refusal loss and its gradient norm. Our detection procedure is shown in Figure 1 (c). Gradient Cuff can be summarized into two steps:
  • (Step 1) Sampling-based Rejection: In the first step, we reject the user query $x$ by checking whether $f_{\theta}(x)<0.5$. If true, then $x$ is rejected, otherwise, $x$ is pushed into Step 2.
  • (Step 2) Gradient Norm Rejection: In the second step, we regard $x$ as having jailbreak attempts if the norm of the estimated gradient $g_{\theta}(x)$ is larger than a configurable threshold $t$, i.e., $\|g_{\theta}(x)\|>t$.
@@
We summarize our method in Algorithm 1. The algorithm is implemented by querying the LLM $T_{\theta}$ multiple times, each to generate a response for the same input query $x$.
The total query times to $T_{\theta}$ required to compute $f_{\theta}(x)$ and $g_{\theta}(x)$ in Gradient Cuff is at most $q=N\cdot(P+1)$. To maintain the LLM’s efficiency, we also explored the use of batch inference to compute these queries in parallel, thereby reducing the total running cost of the LLM. For example, the running time can only be increased by 1.3$\times$ when the total query times were 10$\times$ of the original. See detailed discussion in Appendix A.15.